Reduce each of the following equations into canonical form and find the general solution: (a) Uz + xy = u, (b) ux + x +y = y, (C) ux + 2xy uy = x, (d) Uz – yuy – u = 1.

1
Expert's answer
2021-11-16T16:56:24-0500

c)

"p+2xyq=x"


"\frac{dx}{1}=\frac{dy}{2xy}=\frac{dz}{x}"


"2xdx=dy\/y"


"x^2=lny+c_1"


"xdx=dz"


"x^2=2z+c_2"


"F(c_1,c_2)=F(x^2-lny,x^2-2z)=0"


b)

"\frac{dx}{1}=-\frac{du}{x}=\frac{dy}{0}"

"xdx=-du"


"x^2=-2u+c_1"


"x^2+2u=c_1"


"y=c_2"


"F(c_1,c_2)=F(x^2+2u,y)=0"


d)

"\frac{dz}{1}=-\frac{dy}{y}=\frac{du}{u+1}=\frac{dx}{0}"


"z=-lny+lnc_1"

"e^zy=c_1"


"-lny=ln(u+1)-lnc_2"

"c_2=y(u+1)"


"x=c_3"


"F(c_1,c_2,c_3)=F(e^zy,y(u+1),x)=0"


a)

"\frac{dz}{1}=\frac{du}{u-xy}=\frac{dx}{0}=\frac{dy}{0}"


"x=c_1"

"y=c_2"


"dz=du\/(u-c_1c_2)"


"z=ln(u-c_1c_2)+lnc_3"

"e^z=c_3(u-c_1c_2)=c_3(u-xy)"

"e^z\/(u-xy)=c_3"


"F(c_1,c_2,c_3)=F(x,y,e^z\/(u-xy))=0"